OpenStudy (anonymous):

Y=x^2-6x-8 Find roots by factoring or quadratic formula if factoring doesnt work.

OpenStudy (anonymous):

You sure it's -8?

OpenStudy (anonymous):

yup, y=x^2-6x-8. Yup. Positive

OpenStudy (anonymous):

You're gonna want to use the quadratic formula for this one.

OpenStudy (anonymous):

Thought so. I dunno why, but the first time i tried, i got -38,0 and like -26,0 or something like that. i knew it was completely wrong tho...Haha

OpenStudy (anonymous):

Cause the only integer factors of 8 are 4 and 2, but they can't both be negative (to give -6) without the 8 being positive.

OpenStudy (anonymous):

Hmm..

OpenStudy (anonymous):

So you're gonna get non-integers

OpenStudy (anonymous):

So Decimals... Hopefully theyre only .5

OpenStudy (anonymous):

polpack will you help me?

OpenStudy (anonymous):

K i'll solve it.

OpenStudy (anonymous):

I just leave em as fractions usually

OpenStudy (anonymous):

but then my teachers rarely let us use calculators

OpenStudy (anonymous):

She's fine with us using them. she just likes us to try without them. you can play cards on our calculators..haha

OpenStudy (anonymous):

polpak. can you help me

OpenStudy (anonymous):

\[x=\frac{6 \pm\sqrt{36 - 4(-8)}}{2} = 3 \pm \frac{\sqrt{68}}{2} = 3 \pm \frac{\sqrt{4\times 17}}{2} = 3 \pm \sqrt{17}\]

OpenStudy (anonymous):

but theres no variables in that. :(

OpenStudy (anonymous):

those are the roots, sorry.

OpenStudy (anonymous):

to factor just say \(y = (x-3+\sqrt{17})(x-3-\sqrt{17})\)

OpenStudy (anonymous):

polpak i need help please!

OpenStudy (anonymous):

ohh, so all i need to do is...ohh, okay...THANKS!

OpenStudy (anonymous):

ohh, so all i need to do is...ohh, okay...THANKS!